OpenStudy (anonymous):

1. Solve for x in the following proportion: 3/7 = 2/x A. x = 14 B. x = 7 C. x = 21/2 D. x = 14/3 2. A map is scaled so that 1 cm on the map is equal to 15 actual miles. If two cities on the map are 10 cm apart, proportion would you use to solve the problem? A. map/actual = 1/15 = 10/x

OpenStudy (anonymous):

d answer of first one

OpenStudy (anonymous):

B. map/actual = 15/1 = 10/x C. map/actual = 15/1 = x/10 D. map/actual = x/15 = 10/1

OpenStudy (anonymous):

its again cross multiply

OpenStudy (anonymous):

answers for 2 .

OpenStudy (anonymous):

what do i cross multiply though ? are you talking about the 2nd one ?

OpenStudy (anonymous):

no the first one is about cross multiply

OpenStudy (anonymous):

is there any value given for x

OpenStudy (anonymous):

ohhhh . i see ! okay i was really confused on that .

OpenStudy (anonymous):

no there isn't any /:

OpenStudy (anonymous):

i think it should be d

OpenStudy (anonymous):

alright . i appreciate you for trying to help me out though ! (:

OpenStudy (anonymous):

@alainabbyboo22 its d i am sure look if 1 cm is 15 miles then 10 cm is 150 miles so when we solve all the given equations we see the correct answer is only occurring in d there fore its D. :)

OpenStudy (anonymous):

you're awesome , just throwing that out there (: i seriously can't thank you enough for the help !

OpenStudy (anonymous):

don't bother open study is meant for helping others :) and i also love toooo

OpenStudy (anonymous):

(: